# Post-detection inference for sequential changepoint localization

### v4

@@ H.3 Experiments with detection algorithms that control PFA for Gaussian mean change scenario
100 $\{N(\mu,1):\mu\leq 0.25\}$ $\{N(\mu,1):\mu>0.75\}$ 0.97 0.97 26.02 3.97 125.64
100 $\{N(\mu,1):\mu\leq 0.1\}$ $\{N(\mu,1):\mu>0.9\}$ 0.97 0.97 21.96 3.88 121.28
500 $\{N(\mu,1):\mu\leq 0.25\}$ $\{N(\mu,1):\mu>0.75\}$ 0.97 0.97 25.18 3.82 544.91
500 $\{N(\mu,1):\mu\leq 0.1\}$ $\{N(\mu,1):\mu>0.9\}$ 0.98 0.98 21.45 3.75 536.76
H.3.4 Comparison of experimental results with ARL and PFA controlling detection algorithms
As expected, algorithms that control PFA exhibit a significantly higher average detection delay compared to those controlling ARL. This delay arises because stricter false alarm constraints force the detection algorithm to accumulate stronger evidence before signaling a change.
@@ H.3.4 Comparison of experimental results with ARL and PFA controlling detection algorithms
In terms of inference, we observe that PFA-controlling algorithms often tend to yield slightly lower conditional coverage probabilities with confidence sets of comparable sizes. Also, conditional and unconditional coverage rates are identical when PFA-controlling detection algorithms are employed, due to the very low probability of false detections, while the ARL controlling algorithms lead to lower unconditional conditional coverage rates for $T=500$, supporting our theory. Moreover, we observe that the average absolute deviations of the point estimates are barely smaller with the PFA controlling algorithms.
Overall, the major differences between these two types lie in detection delay and unconditional coverage properties. We do not observe significant differences in the size and conditional coverage properties of the confidence sets nor in the absolute deviation of the point estimates, across these two types of detection algorithms.
H.4 Experimental results with Poisson distribution with detection algorithms that control ARL
H.4.1 Known pre- and post-change (Setting I)
We consider Poisson the mean-change problem: $\text{Pois}(1)$ (pre-change) vs $\text{Pois}(2)$ (post-change) with the true (unknown) changepoint at $T=100,500$. We employ the likelihood-based CUSUM change detection method with thresholds at $A=1000,10000$. After detection, we construct confidence sets (adaptive, Equation 4.1) for $T$. Fig. H.6 provides visualizations of the confidence sets and point estimates of the changepoints across $5$ runs.
Figure H.6: Setting I (known pre- and post-change Poisson distributions): The first T1T-1 observations are drawn from Pois(1)(1) and the rest from Pois(2)(2), with $T=100,500$. The point estimate is shown in a vertical red dashed line, and the confidence set (using Algorithm 2) is shown in red points. $B=N=100$, $\alpha=0.05,L=\infty$. Results of $5$ independent simulations are shown.
We report the average size and the conditional and unconditional coverage rates across $500$ independent runs in Table H.7.
@@ H.4.1 Known pre- and post-change (Setting I)
Table H.7: Setting I: Average (of $500$ independent runs) size and coverage of the confidence sets using Algorithm 2 with $F_{0}=\text{Pois}(1),F_{1}=\text{Pois}(2),B=N=100$, $\alpha=0.05$, the average absolute deviation $\hat{T}$ and average detection delay (given a true detection) with threshold AA, while varying $T$ and $A$.
$T$ AA
Conditional
coverage
Unconditional
coverage
Size $\mathbb{E}_{T}(|\hat{T}-T||\tau\geq T)$ $\mathbb{E}_{T}(\tau-T|\tau\geq T)$
100 10000 0.94 0.94 18.04 4.02 22.15
100 1000 0.94 0.94 17.45 3.98 15.72
500 10000 0.95 0.95 18.53 4.30 22.10
500 1000 0.96 0.89 18.72 3.96 16.14
For $A=10000$, conditional and unconditional coverages are equal, as the probability of false detection before the true change is negligible. However, for $T=500$ with $A=1000$, the presence of a significant number of false detections before the true change results in lower unconditional coverage. Nevertheless, the empirical conditional coverage rates meet the desired level, supporting our theory. We also report the average absolute deviation of $\hat{T}$ and delay (given no false detection) in the table.
H.4.2 Known pre-change and composite post-change (Setting II)
We perform experiments for Poisson rate-change scenarios having pre-change distribution as $\text{Pois}(1)$, and the true (unknown) post-change data distribution as $\text{Pois}(2)$, with the changepoint at $T=100,200,$ assuming the composite post-change model to be $\mathcal{P}_{1}=\{\text{Pois}(\theta):\theta>1.9\}$. For detection, we employ a weighted CUSUM detector defined as
$\tau_{\text{wcusum}}=\inf\left\{\tau\in\mathbb{N}:\max_{1\leq j\leq\tau}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{\tau}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}dW(\theta)\geq A\right\},$ (S5)
for some weight $W(\theta)$ such that $\int_{\theta\in\Theta_{1}}dW(\theta)=1$. We use thresholds $A=1000,10000$ and a discrete weight distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$, $d=0.2$, and $\theta=1.9$, with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$, which appeared to us to be a sensible discretized exponential mixture over the alternative class. After detection, we construct confidence sets (as defined in Section C.2), with $\alpha=0.05,\beta=0.025,B=N=100,L=\infty$.
We report the average size and the conditional and unconditional coverage across $500$ independent runs in (H.8). The empirical conditional coverage rates are always a bit higher than the theoretical bound $1-\alpha-\beta$ (which is 0.925 in our experiments). It is worth noting that both the lengths and the conditional coverage rates are higher compared to the known pre- and post-change settings (Table H.7). We also report the average absolute deviation of $\hat{T}$ and delay (given no false detection) in the table. We observe that the confidence sets in this setting tend to be a bit more conservative compared to those with known pre- and post-change distributions (Setting 1), exhibiting both higher coverage and larger size.
Table H.8: Setting II: Average (of $500$ independent runs) size and (conditional and unconditional) coverage rates of confidence sets Section 3.2. with $B=N=100$ and $\alpha=0.05,\beta=0.025,L=\infty$, the average absolute deviation $\hat{T}$ and average detection delay (given no false detection) of the change detection method used (S11) with threshold AA, having pre- and post-change data-generating distributions as $\text{Pois}(1)$ and $\text{Pois}(2)$ respectively, while we vary the changepoint $T$ and $A$.
@@ H.4.2 Known pre-change and composite post-change (Setting II)
Unconditional
coverage
Size $\mathbb{E}_{T}(|\hat{T}-T||\tau\geq T)$ $\mathbb{E}_{T}(\tau-T|\tau\geq T)$
100 10000 0.97 0.97 33.13 6.11 27.68
100 1000 0.97 0.96 33.85 5.59 19.14
200 10000 0.98 0.98 36.08 5.98 26.81
200 1000 0.97 0.93 35.24 5.74 21.27
H.4.3 Composite pre- and post-change (Setting III)
Now we conduct experiments for Gaussian mean-change scenarios with composite pre- and post-change models having the changepoint at $T=100$ and $500$, the true (unknown) pre- and post-change data distributions as $N(0,1)$ and $N(1,1)$ respectively, assuming the composite pre and post-change models to be $\mathcal{P}_{0}=\{\text{Pois}(\theta):\theta<0.9\}$ and $\mathcal{P}_{1}=\{\text{Pois}(\theta):\theta>1.9\}$ respectively. For change detection, we employ a weighted CUSUM-type detector, where the “closest” element from the pre-change (the Reverse Information Projection (RIPr) of the post-change element to the pre-change class) is used as a representative of the pre-change model, which is defined as
$\tau_{\text{wcs-ripr}}=\inf\left\{\tau\in\mathbb{N}:\max_{1\leq j\leq\tau}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{\tau}\frac{f_{\theta}(X_{i})}{f_{\theta}^{*}(X_{i})}dW(\theta)\geq A\right\},$ (S6)
where $f_{\theta}^{*}$ is the RIPr of $f_{\theta}$ on $\mathcal{P}_{0}$. For the post-change parameter, the same discrete weight distribution as in Section H.4.2 is considered.
After detection, we construct confidence sets (as defined in Section D.2, assuming $T\neq 1$), with $\alpha=0.05,\beta=\gamma=0.025,B=N=100,L=\infty$. We report the average size, conditional and unconditional coverage rates, average absolute deviation of $\hat{T}$, and delay (given no false detection) across $500$ independent runs in Table H.9.
Table H.9: Setting III: Average (of $500$ independent runs) conditional and unconditional coverage and size of confidence sets with $N=B=100$ and $\alpha=0.05,\beta=\gamma=0.025$,the average absolute deviation $\hat{T}$ and average detection delay (given no false detection) of the change detection method with threshold AA, having pre- and post-change data-generating distributions as $\text{Pois}(1)$ and $\text{Pois}(2)$ respectively, while we vary the changepoint $T$ and $A$.
T A
Conditional
coverage
Unconditional
coverage
Size $\mathbb{E}_{T}(|\hat{T}-T||\tau\geq T)$ $\mathbb{E}_{T}(\tau-T|\tau\geq T)$
100 10000 0.97 0.97 35.04 6.34 31.56
100 1000 0.97 0.97 35.19 6.27 26.33
200 10000 0.98 0.98 37.54 5.92 30.45
200 1000 0.98 0.97 37.86 6.18 25.82
However, the empirical conditional coverage rates are always a bit higher than the theoretical bound $1-\alpha-\beta-\gamma$ (which is 0.9 in our experiments). It is worth noting that both the lengths and the conditional coverage rates are higher compared to the known pre- and post-change settings (Table H.7). We observe that the confidence sets in this setting tend to be a bit more conservative compared to those with known pre- and post-change distributions (Table H.7), exhibiting both higher coverage and larger size.
Appendix I A heuristic method for composite pre- or post-change
This section presents an alternative method for constructing confidence intervals for composite pre- and post-change distributions. Specifically, we propose a heuristic bootstrap-based approach that estimates the unknown distributions from the observed data and subsequently generates simulations from the estimated distributions.
I.1 Known pre-change and unknown post-change distributions
Here we assume that the pre-change distribution $F_{0}$ is known and the composite post-change model is $\mathcal{P}_{1}$. Building on the approach from previous sections, we define the null hypothesis that a changepoint occurs at time $t\in\{1,\cdots,\tau\}$ as:
$\mathcal{H}_{0,t}:X_{1},\cdots,X_{t-1}\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{0}\text{ and }X_{t},X_{t+1},\cdots\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{1},\text{ for some (unknown) }F_{1}\in\mathcal{P}_{1}.$
For testing the null hypothesis, we define some test statistic $M_{t}$, based on $X_{1},\cdots,X_{\tau}$. Since the pre-change $F_{0}$ is known, we can generate $N$ many i.i.d. sequences under $F_{0}$ until a change is detected, recording the detection times as $\tau_{1},\cdots,\tau_{N}$, for some $N\in\mathbb{N}$. Now, the challenge in conducting a level $\alpha\times\mathbb{\mathbb{P}}_{\infty}(\tau\geq t)$ test for each $t$ lies in the composite nature of $\mathcal{H}_{0,t}$, which prevents us from straightforwardly simulating data to determine a rejection threshold as we did in Algorithm 1. To address this, we employ a simple bootstrap-based approach. The main idea is to first obtain some point estimate of the changepoint (for example, a likelihood ratio based estimate, as defined earlier). We denote the estimate as $\hat{T}$. Then we use the observations from $\hat{T}$ to $\tau$ to estimate the post-change distribution $F_{1}$, denoted $\hat{F}_{1}$. Then, for some $B\in\mathbb{N}$, we generate $B$ independent bootstrap data streams with known pre-change distribution $F_{0}$ and estimated post-change distribution $\hat{F}_{1}$ having changepoint at $t$.
For each sampled stream $j$, we compute the same test statistic $M_{t}^{j}$ under the null hypothesis, resulting in statistics $M_{t}^{1},M_{t}^{2},\dots,M_{t}^{B}$. We reject $\mathcal{H}_{0,t}$ if
$M_{t}>\operatorname{Quantile}(1-\alpha\times\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N;M_{t},M_{t}^{1},\cdots,M_{t}^{B}).$
Our confidence set is the collection of all $t$ before $\tau$, for which we fail to reject $\mathcal{H}_{0,t}$. Algorithm 4 contains an overview of the method.
Algorithm 4 CI for known pre-change $F_{0}$ and composite post-change $\mathcal{P}_{1}$
[…]
2:A confidence set $\mathcal{C}$ for changepoint $T$
3:Compute a point estimator $\hat{T}$ for $T$
4:Obtain an estimate $\hat{F}_{1}$ of the (unknown) post-change $F_{1}\in\mathcal{P}_{1}$ based on $X_{\hat{T}},\cdots,X_{\tau}$
5:for $i=1$ to NN do
6:  Simulate $X_{1}^{i},X_{2}^{i},\cdots\stackrel{{\scriptstyle iid}}{{\sim}}F_{0}$, until change is detected at $\tau_{i}$ using $\mathcal{A}$, or until $\tau$, whichever happens first
7:end for
[…]
9:for $t=1$ to $\tau$ do
10:  Compute some test statistics $M_{t}$ based on $\tau$ and $X_{1},\cdots,X_{\tau}$
11:  for $j=1$ to $B$ do
12:   Simulate $X_{1}^{j},\cdots,X_{t-1}^{j}\stackrel{{\scriptstyle iid}}{{\sim}}F_{0};\penalty 10000\ X_{t}^{j},X_{t+1}^{j},\cdots\stackrel{{\scriptstyle iid}}{{\sim}}\hat{F}_{1}$, until a change is detected at $\tau_{j}^{\prime}$ using $\mathcal{A}$
13:   Compute the same test statistics $M_{t}^{j}$ based on $\tau_{j}^{\prime}$ and $X_{1}^{j},\cdots,X_{\tau_{j}^{\prime}}^{j}$
14:  end for
15:  if $M_{t}\leq\operatorname{Quantile}(1-\alpha\times\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N;M_{t},M_{t}^{1},\cdots,M_{t}^{B})$ then
16:   $\mathcal{C}=\mathcal{C}\cup\{t\}$
17:  end if
18:end for
I.2 Composite pre- and post-change distributions
We now consider the general case where both pre and post-change distributions follow some composite models $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$ respectively. Following the approach from the previous sections, we define the null hypothesis that a changepoint occurs at time $t\in\{1,\cdots,\tau\}$ as:
[…]
For testing the null hypothesis, we define some test statistic $M_{t}$, based on $X_{1},\cdots,X_{\tau}$ (eg., the likelihood-ratio-based test statistic as defined earier).
As in Algorithm 4, we employ a simple bootstrap-based approach, but here we need to estimate both pre- and post-change distributions. We first obtain some point estimate $\hat{T}$ of the changepoint $T$. Then we use the observations from $1$ to $\hat{T}-1$ to estimate the pre-change distribution $F_{0}$, denoted $\hat{F}_{0}$ and from $\hat{T}$ to $\tau$ to estimate the post-change distribution $F_{1}$, denoted $\hat{F}_{1}$. For some $N\in\mathbb{N}$, we generate $N$ many i.i.d. sequences under $\hat{F}_{0}$ until a change is detected and denote the detection times as $\tau_{1},\cdots,\tau_{N}$. Then, for some $B\in\mathbb{N}$, we generate $B$ independent bootstrap data streams with the estimated pre- and post-change distributions $\hat{F}_{0}$ and $\hat{F}_{1}$ respectively, having changepoint at $t$. For each sampled stream $j$, we calculate the test statistic $M_{t}^{j}$ to obtain $M_{t}^{1},M_{t}^{2},\dots,M_{t}^{B}$. We reject $\mathcal{H}_{0,t}$ if we have
$M_{t}>\operatorname{Quantile}(1-\alpha\times\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N;M_{t},M_{t}^{1},\cdots,M_{t}^{B}).$
Our confidence set is the collection of all $t$ before $\tau$, for which we fail to reject $\mathcal{H}_{0,t}$. Algorithm 5 contains an overview of the method.
Algorithm 5 CI for unknown pre- and post-change
[…]
2:A confidence set $\mathcal{C}$ for changepoint $T$
3:Compute some point estimator $\hat{T}$ for $T$
4:Obtain an estimate $\hat{F}_{0}$ of the (unknown) pre-change $F_{0}\in\mathcal{P}_{0}$ based on $X_{1},\cdots,X_{\hat{T}-1}$
5:Obtain an estimate $\hat{F}_{1}$ of the (unknown) post-change $F_{1}\in\mathcal{P}_{1}$ based on $X_{\hat{T}},\cdots,X_{\tau}$
6:for $i=1$ to NN do
7:  Simulate $X_{1}^{i},X_{2}^{i},\cdots\stackrel{{\scriptstyle iid}}{{\sim}}\hat{F}_{0}$, until change is detected at $\tau_{i}$ using $\mathcal{A}$, or until $\tau$, whichever happens first
8:end for
[…]
10:for $t=1$ to $\tau$ do
11:  Compute some test statistics $M_{t}$ based on $X_{1},\cdots,X_{\tau}$
12:  for $j=1$ to $B$ do
13:   Simulate $X_{1}^{j},\cdots,X_{t-1}^{j}\stackrel{{\scriptstyle iid}}{{\sim}}\hat{F}_{0};\penalty 10000\ X_{t}^{j},X_{t+1}^{j},\cdots\stackrel{{\scriptstyle iid}}{{\sim}}\hat{F}_{1}$, until a change is detected at $\tau_{j}$ using $\mathcal{A}$
14:   Compute the same test statistics $M_{t}^{j}$ based on $X_{1}^{j},\cdots,X_{\tau_{j}}^{j}$
15:  end for
16:  if $M_{t}\leq\operatorname{Quantile}(1-\alpha;M_{t},M_{t}^{1},\cdots,M_{t}^{B})$ then
17:   $\mathcal{C}=\mathcal{C}\cup\{t\}$
18:  end if
19:end for
I.3 Experimental Results
I.3.1 Known pre-change and composite post-change (Setting II)
We conduct experiments for both Normal and Poisson mean-change scenarios with a composite post-change model having the changepoint at $T=100$. For detection, we employ a weighted CUSUM detector (S11) with a threshold at $A=1000$ and a discrete weight distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$ (for normal distribution, we use $\theta=d=0.2$ and for Poisson distribution $\theta=1.2,d=0.2$) with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$, which appeared to us to be a sensible discretized exponential mixture over the alternative class. After detection, we employ Algorithm 4 to construct confidence sets. Table I.1 summarizes the results, including the average size and the conditional and unconditional coverage across $500$ independent runs. For $T=100$, conditional and unconditional coverages are equal, as the probability of false detection before the true change is negligible. However, for $T=500$ with $A=1000$, the presence of a significant number of false detections before the true change results in lower unconditional coverage. Nevertheless, the conditional coverage is maintained at the desired level, even though we do not have a theoretical guarantee in this case.
Table I.1: Setting II: Average (of $500$ independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 4 with $B=100,N=100$ and $\alpha=0.05$, using weighted CUSUM change detection method (S11) with threshold $A=1000$.
T Pre-change
Composite
post-change
True (unknown)
post-change
Conditional
coverage
Unconditional
@@ I.3.1 Known pre-change and composite post-change (Setting II)
100 $N(0,1)$ $\{N(\mu,1):\mu>0\}$ $N(1,1)$ 0.94 0.94 14.76
100 $\text{Pois}(1)$ $\{\text{Pois}(\lambda):\lambda>1\}$ $\text{Pois}(2)$ 0.95 0.95 20.95
500 $N(0,1)$ $\{N(\mu,1):\mu>0\}$ $N(1,1)$ 0.95 0.91 14.06
500 $\text{Pois}(1)$ $\{\text{Pois}(\lambda):\lambda>1\}$ $\text{Pois}(2)$ 0.95 0.89 20.83
I.3.2 Composite pre- and post-change (Setting III)
Now we conduct experiments for both Normal and Poisson mean-change scenarios with composite pre- and post-change models having the changepoint at $T=100$ and $500$. We use a weighted CUSUM detector with RIPr defined in (S12) for change detection. The thresholds are set at $A=1000$ for both $T=100$ and $T=500$. We considered the same discrete weight distribution as above, taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$ (for normal distribution, we use $\theta=0.8,d=0.2$ and for Poisson distribution $\theta=1.8,d=0.2$) with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$, which appeared to us to be a sensible discretized exponential mixture over the alternative class. After detection, we employ Algorithm 5 to construct confidence sets. We report the average size and the conditional and unconditional coverage across $500$ independent runs in Table I.2. The conditional coverage is maintained at the desired level, even though we do not have a theoretical guarantee in this setting.
Table I.2: Setting III: Average (of $500$ independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 5 with $N=B=100$ and $\alpha=0.05$, using weighted CUSUM type change detection method (S12) with threshold $A=1000$.

### v6

@@ F.1.3 Composite pre- and post-change (Setting C)
100 2 0.97 0.97 23.54 50.64
100 3 0.97 0.97 21.82 82.95
500 2 0.97 0.97 21.37 122.84
500 3 0.98 0.98 23.22 193.59
F.2 Comparison of experimental results with ARL and PFA controlling detection algorithms
@@
As expected, algorithms that control PFA exhibit a significantly higher average detection delay compared to those controlling ARL. This delay arises because stricter false alarm constraints force the detection algorithm to accumulate stronger evidence before signaling a change. In terms of inference, we observe that PFA-controlling algorithms often tend to yield slightly lower conditional coverage probabilities with confidence sets of comparable sizes. Also, conditional and unconditional coverage rates are identical when PFA-controlling detection algorithms are employed, due to the very low probability of false detections, while the ARL controlling algorithms lead to lower unconditional conditional coverage rates for larger $T$, ie.e., $T=500$, supporting our theory. Overall, the major differences between these two types lie in detection delay and unconditional coverage properties. We do not observe significant differences in the size and conditional coverage properties of the confidence sets, nor in the absolute deviation of the point estimates, across these two types of detection algorithms.
Appendix G Some additional experimental results
G.1 Experiments varying $T$ and ARL
We have seen in the derivation of the asymptotic length bound that it increases with the ratio $T/A_{T}$. Now, we experimentally validate the fact by conducting the same experiment as in Setting I (known pre- and post-change) in Section 6 with varying $T$ thresholds $A$ of the CUSUM detector (larger threshold means large ARL). Average (of $500$ independent runs) size and coverage of the confidence sets using Equation 10 are reported in Table 14. The results show a clear pattern: as the ratio $T/A$ increases, the resulting confidence sets become wider, reflecting growing conservativeness.
@@ G.1 Experiments varying $T$ and ARL
Table 14: Setting I: Known pre- and post-change. Average (of $500$ independent runs) size and coverage of the confidence sets using Equation 10 with $F_{0}=N(0,1),F_{1}=N(1,1),N=100$, $\alpha=0.1$, the average detection delay (given a true detection) with threshold AA, while varying $T$ and $A$.
@@
Marginal
@@
Conditional
Delay
50 100 0.978 0.91 14.382 9.33
1000 100 0.989 0.21 18.283 9.14
2000 100 0.992 0.07 19.627 8.35
500 1000 0.925 0.97 12.618 13.22
10000 1000 0.947 0.27 18.254 13.25
20000 1000 1 0.07 22.113 13.22
G.2 Experiments with dependent pre- and post-change data
[…]
We use the likelihood ratio-based CUSUM detector:
After detection, we compute the adaptive confidence sets (17) with the test statistic in (9) with $R^{(t)}_{n}=\frac{1}{2}\left(\frac{p_{0}}{p_{1}}\right)^{n_{10}(t:n)}\left(\frac{1-p_{0}}{1-p_{1}}\right)^{n_{11}(t:n)}$, $S^{(t)}_{n}=\frac{1}{2}\left(\frac{p_{1}}{p_{0}}\right)^{n_{10}(n:t-1)}\left(\frac{1-p_{1}}{1-p_{0}}\right)^{n_{11}(n:t-1)}$, $B=N=100$ and $\alpha=0.1$. Note that $\{S^{(t)}_{n}\}_{n<t}$ is not a valid e-process in the backward filtration. In contrast to our universal method, the adaptive procedure in Algorithm 1 (or Equation 17) accommodates arbitrary test statistics and does not rely on e-process structure; consequently, it remains applicable in settings with dependent data. Table 15 contains the results averaged across $100$ random runs.
Table 15: Known pre- and post-change with dependency